OpenStudy (lgbasallote):

i need to practice on finding minimum/maximum can someone give me something and check my answer afterwards? it'll help a lot...

OpenStudy (anonymous):

find the max and min of \(xe^x\) on \([-2,1]\)

OpenStudy (lgbasallote):

hmm okay...so the first derivative of \(xe^x\) i know to be \[xe^x + e^x\] if i equate to 0 \[e^x(x + 1) = 0\] i have x = -1 what do i do with [-2. 1]?

OpenStudy (anonymous):

you have the critical point, to find the max and min, you have to check the critical point and also the endpoint of the interval biggest is max smallest is min

OpenStudy (anonymous):

try \(x^2\ln(x)\) for a global min

OpenStudy (lgbasallote):

how do i "check the critical point and also the endpoint of the interval" substitute?

OpenStudy (anonymous):

or for an easier one minimum value of \(7x+\frac{48}{x}\) on \((0,\infty)\)

OpenStudy (anonymous):

yes substitute

OpenStudy (anonymous):

you need to check both the critical points and also the endpoints of the interval

OpenStudy (lgbasallote):

where do i substitute? in the original equation? like in xe^x i sub in x = -1 so -1/e?

OpenStudy (anonymous):

yes that is the minimum

OpenStudy (anonymous):

don't forget the minimum value is the (y\) value, not the \(x\) value calculus find the \(x\) value to give the max and min, but min in the output not the input

OpenStudy (anonymous):

try \[\frac{x^2}{|x-2|}\] on \((2,5)\) for a challenge

OpenStudy (lgbasallote):

so what was the maximum for that xe^x?

OpenStudy (lgbasallote):

wait...just to clarify that [-2,1] was just to test if x = -1 is within the given limits right?

OpenStudy (anonymous):

you have to check endpoints as well at \(-2\) you get \(-2e^{-2}=-\frac{2}{e^2}\) at \(-1\) you get \(-e^{-1}=-\frac{1}{e}\) and at 1 you get \(e\)

OpenStudy (anonymous):

compare the numbers you see that \(-\frac{1}{e}\) is the smallest at \(e\) is the largest

OpenStudy (anonymous):

if you want more, let me know

OpenStudy (lgbasallote):

ohhh so e is the max and -1/e is the min?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it is on the interval \([-2,1]\)

OpenStudy (lgbasallote):

ahhh nice!!

OpenStudy (lgbasallote):

what does global min mean?

OpenStudy (anonymous):

minimum value over the entire domain of the function

OpenStudy (lgbasallote):

ohh so no limits...i assume that's where second derivative comes in?

OpenStudy (anonymous):

no you can use the first derivative if you like

OpenStudy (anonymous):

or second to test whether you have a max or min

OpenStudy (lgbasallote):

\[x^2 \ln x\] \[y' = x + 2x \ln x\] \[x(1 + 2 \ln x = 0\] \[x = 0 \; \ln x = -\frac 12 \implies x = e^{-1/2}\] i would say e^-1/2 is the global min yes?

OpenStudy (lgbasallote):

\[x = 0 \; ; \; \ln x = -\frac 12 \implies x = e^{-1/2}\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

and the reason that \(e^{-\frac{1}{2}}\) is not the global minimum is that it is the \(x\) value

OpenStudy (anonymous):

don't forget, the minimum is not the input, it is the output

OpenStudy (lgbasallote):

ohh substitute...

OpenStudy (anonymous):

right!

OpenStudy (lgbasallote):

\[e^{-1/2 \times 2} \ln (e^{-1/2}) \implies e^{-1} \times -\frac 12 \implies -\frac{1}{2e}\] right?